Find the integral surface of the equation pq=z passing through C : x0 = 0, y0 = s, z0 = s2

1
Expert's answer
2021-02-24T07:40:52-0500

We use the Charpit's method for solution.

"f(x,y,z,p,q)=pq-z=0" with "p=z_x" and "q=z_y".

Charpit's auxiliary equations in general case:

"\frac{-dx}{\frac{\partial f}{\partial p}}= \frac{-dy}{\frac{\partial f}{\partial q}}= \frac{-dz}{p\frac{\partial f}{\partial p} +q\frac{\partial f}{\partial q}}=\frac{dp}{\frac{\partial f}{\partial x}+p\frac{\partial f}{\partial z}}=\frac{dq}{\frac{\partial f}{\partial y}+q\frac{\partial f}{\partial z}}"

Charpit's auxiliary equations for the given PDE:

"\frac{-dx}{q}= \frac{-dy}{p}= \frac{-dz}{2pq}=\frac{dp}{-p}=\frac{dq}{-q}"

One of the possible solutions is that with pq=0. In this case z=pq=0, "p=z_x=0" and "q=z_y=0" .

We assume from now that "pq\ne0" .

"\frac{-dx}{q}=\frac{dq}{-q}" implies that "dq=dx", "q=x+a"

"\frac{-dy}{p}= \frac{dp}{-p}" implies that "dp=dy", "p=y+b"

Finally, z=pq=(x+a)(y+b).

The integral surface passing through C : x0 = 0, y0 = s, z0 = s2 must satisfy:

"s^2=a(b+s)". Therefore, "a\ne 0" and "b=a^{-1}s^2-s".

Answer. "z=(x+a)(y+a^{-1}s^2-s)"